# MultiGripperGrasp: A Dataset for Robotic Grasping from Parallel Jaw Grippers to Dexterous Hands

### v1

@@ Abstract
We introduce a large-scale dataset named MultiGripperGrasp for robotic grasping. Our dataset contains 30.4M grasps from 11 grippers for 345 objects. These grippers range from two-finger grippers to five-finger grippers, including a human hand. All grasps in the dataset are verified in Isaac Sim to classify them as successful and unsuccessful grasps. Additionally, the object fall-off time for each grasp is recorded as a grasp quality measurement. Furthermore, the grippers in our dataset are aligned according to the orientation and position of their palms, allowing us to transfer grasps from one gripper to another. The grasp transfer significantly increases the number of successful grasps for each gripper in the dataset. Our dataset is useful to study generalized grasp planning and grasp transfer across different grippers.  [^1] Data, code and videos for the project are available at
@@ III-A Objects and Grippers
Object Set. We chose a subset of 329 objects from the GoogleScannedObjects dataset [36] as our core object set. These are 3D scanned objects in everyday settings. The subset is chosen according to the FewSOL [37] dataset for few-shot object recognition in robotic settings. Furthermore, we also included 16 objects from the YCB object set [38] due to the popularity of the YCB objects in robot manipulation. Some examples of objects in our dataset are shown in Fig. 3. We do not have access to the ground truth mass and friction coefficient of these objects. Therefore, when importing these objects into Isaac Sim, we used a fixed density (100 $kg/m^{3}$) to compute object mass and used a friction coefficient of 0.5.
@@
For each gripper, we manually selected preferred contact regions on the gripper surface (fingers and palm), which is common in contact-based grasp optimization methods [26, 6]. To transfer grasps between different grippers, a common notion of gripper pose was used as illustrated by the axes in Fig. 4. The translation refers to the palm center of the gripper, and the orientation is with respect to a canonical pose of the gripper palm pointing in a fixed direction. Thus, any grasp pose from a gripper can be transferred to another by using this pose alignment. For collision checking of these grippers in Isaac Sim, we use a convex decomposition over the gripper meshes.
Figure 5: Illustration of the grasp ranking process with Isaac Sim.
@@ III-C Grasp Labeling
Motivation. To further evaluate the quality of the grasps generated by GraspIt!, we use the Isaac Sim [20] robotics simulator to verify these grasps, since GraspIt! only considers object and gripper geometry for grasp analysis. Crucially, we provide a ranked order for these grasps based on the simulation statistics instead of simply providing grasps that pass the simulation test as positive grasps. This allows some grasps in the dataset to be used as negative examples in learning-based methods that use grasp evaluators such as 6D GraspNet [5]. These negative grasps are of higher quality than random sampled grasps, since they represent hard negative examples during training.
Procedure. The general idea of grasp labeling is to check if a grasped object will fall or not due to gravity in simulation as illustrated in Fig. 5. For every grasp generated from GraspIt!, we load the corresponding object and gripper into Isaac Sim. The gripper is placed so that its palm normal is aligned with the direction of gravity, and the object is initialized according to its pose with respect to the gripper. The joint positions of the gripper fingers are initialized according to the given grasp. Then we designed a controller to gradually close the gripper fingers until contact with the object is established. Contact between the fingers and the object triggers the gravity force. After that, physics is simulated over a fixed time interval or until the object falls. The grasp test time was set to 3 seconds.
Stability Measure. Grasps are evaluated according to the time stamp when the object falls. When gravity is applied, a timer is started. Then we recorded the fall-off time for the object as a grasp quality measurement, with higher values for more stable grasps. The object fall-off time is recorded either when the object loses contact with the gripper fingers or when the total test time has expired. In the latter case, the grasp will be given the total test time as the fall-off time. Furthermore, the object is considered to have fallen when the current pose is 0.5 meters away from its initial position. If the gripper cannot make contact with the object, a negative value is assigned as the fall-off time, and the grasp is considered as a failed grasp. The gripper contacts with the object were calculated using the Rigid Body API from Isaac Sim [20]. Fig. 6 shows some ranked grasps according to their object fall-off time.
@@ III-D Controller Design
We need to design a controller to perform the closing motion for the grippers. For parallel jaw grippers, the control can be easily performed by setting the two fingers in their closing positions. For multi-finger grippers, the controller design is not straightforward. We cannot simply change all the finger joints since there are some joints that must remain static, e.g., to avoid lateral movement of the fingers. In this work, we manually identify a subset of joints and their joint values for the closing configuration for each gripper and use a position controller to control these joints to reach the closing configuration. We implemented the following two versions of the controller.
Position Control for Grasp Ranking. For grasps generated from GraspIt!, most of them are already in contact with the object. In this case, we simply control the root joints of the fingers to reach their closing positions (see Fig. 7(a)). The closing positions are manually specified for each gripper.
Position Control for Grasp Transfer. To transfer a grasp from a source gripper to a target gripper, the joint configuration of the source gripper cannot easily be mapped to the target gripper. Therefore, we first set the joint configuration of the target gripper to the open configuration. Then the controller closes all fingers until it comes into contact with the object. After that, the controller set the root joints of the fingers to reach their closing position (see Fig. 7(b)).
@@ IV-A Grasp Ranking
From Table II, we can see that the Robotiq gripper with 3 fingers achieves the highest success rate. We observed low success rates for three grippers in our dataset: the H5 Hand, the Shadow Hand, and the human hand. For the H5 hand, we observed that GraspIt! had difficulties in planning good grasps due to the presence of mimic joints that are not supported. Therefore, the grasps through simulated annealing were sub-optimal. In the case of Shadow hand and human hand, we observed that our position controller is sub-optimal for many grasps when considering highly articulated grippers with high number of DOFs. The question of how to design a better controller for multi-finger hands remains a challenging task.
@@ IV-B Grasp Transfer
We utilize the alignment between grippers as described in Sec. III-A to transfer grasps. We first transform a source gripper pose to its aligned pose, and then transform the aligned pose to the target gripper. Fig. 8 illustrates an example for grasp transfer. To execute the new grasp, the target gripper is fully opened first and then closed using the position controller described in Sec. III-D. The controller performs a closing motion. The joint values at the moment of contact between the gripper and object are recorded, and the object fall-off time is simulated.
Table III shows the confusion matrix of grasp transfer in our experiment. To limit the runtime, we sampled 500 successful grasps per object-gripper pair, where a successful grasp is a grasp with an object fall-off time greater than three seconds. The table shows that many successful grasps of one gripper can be transferred to other grippers. For example, for the Franka Panda gripper, 561K transferred grasps were evaluated as successful by the simulation. For the H5 Hand, where the original grasps generated from GraspIt! are poor, we were able to generate 598K new successful grasps through grasp transfer compared to the 24K original successful grasps.
@@ VI CONCLUSION
In this work, we present MultiGripperGrasp, a large-scale grasping dataset with 11 different grippers across the spectrum of two to five finger grippers. Moreover, we go beyond binary labels of success or failure for grasps and include a ranked order based on grasping performance in a physics simulation along with aligning the gripper poses for grasp transfer. Incorporating different classes of grippers also bring out interesting observations in our experiments, especially with respect to leveraging grasp transfer as a simple baseline for augmenting a gripper with lower performing grasps. With MultiGripperGrasp, we hope to help research on generalized grasping [8, 9] and allow the learning of cross-functional skills between highly different gripper morphologies.
Acknowledgement.

### v2

@@ Abstract
We introduce a large-scale dataset named MultiGripperGrasp for robotic grasping. Our dataset contains 30.4M grasps from 11 grippers for 345 objects. These grippers range from two-finger grippers to five-finger grippers, including a human hand. All grasps in the dataset are verified in the robot simulator Isaac Sim to classify them as successful and unsuccessful grasps. Additionally, the object fall-off time for each grasp is recorded as a grasp quality measurement. Furthermore, the grippers in our dataset are aligned according to the orientation and position of their palms, allowing us to transfer grasps from one gripper to another. The grasp transfer significantly increases the number of successful grasps for each gripper in the dataset. Our dataset is useful to study generalized grasp planning and grasp transfer across different grippers.  [^1] Data, code and videos for the project are available at
@@ III-A Objects and Grippers
Object Set. We chose a subset of 329 objects from the GoogleScannedObjects dataset [36] as our core object set. These are 3D scanned objects in everyday settings. The subset is chosen according to the FewSOL [37] dataset for few-shot object recognition in robotic settings. Furthermore, we also included 16 objects from the YCB object set [38] due to the popularity of the YCB objects in robot manipulation. Some examples of objects in our dataset are shown in Fig. 3. The ground truth mass and friction coefficients of each object is unknown to us, consequently, we used a fixed density (100 $kg/m^{3}$) and friction coefficients (0.5).
Figure 5: Illustration of the grasp ranking process with Isaac Sim.
@@
Methods using classical grasp quality metrics [26, 6] and contact-energy optimization require specifying contact regions on the gripper fingers and palm. For each gripper, we selected such regions using the graphical interface in Graspit. To transfer grasps between different grippers, a common notion of gripper pose was used as illustrated by the axes in Fig. 4. The translation refers to the palm center of the gripper, and the orientation is with respect to a canonical pose of the gripper palm pointing in a fixed direction. Thus, any grasp pose from a gripper can be transferred to another by using this pose alignment. For collision checking of these grippers in Isaac Sim, we use a convex decomposition over the gripper meshes.
@@ III-C Grasp Labeling
Motivation. To further evaluate the quality of the grasps generated by GraspIt!, we use a robotics simulator known as Isaac Sim [20] to verify these grasps, since GraspIt! only considers object and gripper geometry for grasp analysis. Crucially, we provide a ranked order for these grasps based on the simulation statistics instead of simply providing grasps that pass the simulation test as positive grasps. This allows some grasps in the dataset to be used as negative examples in learning-based methods that use grasp evaluators such as 6D GraspNet [5]. These negative grasps are of higher quality than random sampled grasps, since they represent hard negative examples during training.
Procedure. The general idea of grasp labeling is to check if a grasped object will fall or not due to gravity in simulation as illustrated in Fig. 5. To test the grasps in Isaac Sim, the gripper is placed so that its palm normal is aligned with the direction of gravity, and the object is initialized according to its pose with respect to the gripper. This set up allowed us to reliably make gravity a adversarial force to every grasp. The joint positions of the gripper fingers are then initialized according to the given grasp. We designed a controller to gradually close the gripper fingers until contact with the object is established. Contact between the fingers and the object triggers the gravity force. Subsequently, physics is simulated over a fixed time interval or until the object falls. The maximum grasp test time was set to 3 seconds.
Stability Measure. Grasps are evaluated according to the time stamp when the object falls. When gravity is applied, a timer is started. Then we recorded the fall-off time for the object as a grasp quality measurement, with higher values for more stable grasps. The object fall-off time is recorded either when the object loses contact with the gripper fingers or when the maximum test time has expired. In the latter case, the grasp will be given the maximum test time as the fall-off time. Furthermore, the object is considered to have fallen when the current pose is 0.5 meters away from its initial position. If the gripper cannot make contact with the object, a negative value is assigned as the fall-off time, and the grasp is considered as a failed grasp. The gripper contacts with the object were calculated using the Rigid Body API from Isaac Sim [20]. Fig. 6 shows some ranked grasps according to their object fall-off time.
@@ III-D Controller Design
The grasps, when initialized within Isaac Sim, are in contact or close to contact with the objects. However, a gripper controller is needed to exert force on the object to perform the grasp. For parallel jaw grippers, the control can be performed by setting the two fingers to their close position. On the other hand, for multi-finger grippers some joints must remain static, e.g., lateral movement of the fingers. Given the large amount of gripper, object, and pose combinations within our dataset, we opted for a simple controller that could be easily used with all the grippers. For each gripper, we manually identify a subset of joints to move and use a position controller to reach the gripper closing configuration. The following versions of the controller were implemented:
Position Control for Grasp Ranking. For grasps generated from GraspIt!, We simply control the root joints of the fingers to reach their closing positions (see Fig. 7(a)). The closing positions are manually specified for each gripper.
Position Control for Grasp Transfer. To transfer a grasp from a source gripper to a target gripper, we first set the joint configuration of the target gripper to the open configuration. Then the controller closes all fingers until it comes into contact with the object. After that, the controller sets only the root joints of the fingers to reach their closing position (see Fig. 7(b).
@@ IV-A Grasp Ranking
From Table II, we can see that the Robotiq gripper achieves the highest success rate. We observed low success rates for three grippers in our dataset: the H5 Hand, the Shadow Hand, and the human hand. For the H5 hand, GraspIt! had difficulties in planning good grasps due to the presence of unsupported mimic joints. Therefore, the grasps through simulated annealing were sub-optimal. In the case of Shadow hand and human hand, we observed that our controller is sub-optimal for many grasps when considering highly articulated grippers with complex joints. We observed that the generated joint values for the thumbs made the optimal control non-constant. In some cases, the joints may be aligned in a way that our controller creates a stable grasp, while in others it applies adversarial forces to the grasp. The question of how to design a better controller for multi-finger hands remains a challenging task.
@@ IV-B Grasp Transfer
We utilize the alignment between grippers as described in Sec. III-A to transfer grasps. We first transform a source gripper pose to its aligned pose, and then transform the aligned pose to the target gripper. Fig. 8 illustrates an example for grasp transfer. To execute the new grasp, the target gripper is controlled using the position controller described in Sec. III-D. The joint values at the moment of contact between the gripper and object are recorded, and the object fall-off time is simulated.
Table III shows the confusion matrix of grasp transfer in our experiment. To limit the runtime, we sampled 500 successful grasps per object-gripper pair, where a successful grasp is a grasp with an object fall-off time greater than three seconds. Many successful grasps of one gripper were successful grasps in other grippers when transferred. For example, for the Franka Panda gripper, 561K transferred grasps were evaluated as successful by the simulation. For the H5 Hand, where the original grasps generated from GraspIt! are poor, we were able to generate 598K new successful grasps from other grippers compared to the 24K original successful grasps. The table also shows that some grasps from similar grippers were not successful. Gripper reach and morphology are not taken into account when transferring grasps, thus some grasps may be bad when tested in grippers with different finger shape and size.
@@ VI CONCLUSION
In this work, we present MultiGripperGrasp, a large-scale grasping dataset with 11 different grippers across the spectrum of two to five finger grippers. Moreover, we go beyond binary labels of success or failure for grasps and include a ranked order based on grasping performance in a physics simulation along with aligning the gripper poses for grasp transfer. Incorporating different classes of grippers also brings out interesting observations in our experiments, especially with respect to leveraging grasp transfer as a simple baseline for augmenting a gripper with lower performing grasps. With MultiGripperGrasp, we hope to help research on generalized grasping [8, 9] and allow the learning of cross-functional skills between highly different gripper morphologies.
Acknowledgement